# Hook-Lengths, Symplectic/Orthogonal Contents and Amdeberhan’s Conjectures

### v2

@@ 4.1. Partition function of double symplectic Schur functions
In this subsection, we obtain a formula for the partition function consisting of double symplectic Schur functions as in Theorem 4.2. We first need the following commutation relation.
@@
Use the symmetric variables $(\bm{y})=(y_{i})$ and write $p_{k}$ as $p_{k}(\bm{y})$, the $k$-th power sum of the variables $(\bm{y})$. Then, Theorem 4.2 is equivalent to the following.
Corollary 4.3.
We have
@@
Proof: This is similar to the proof of Corollary 3.2. One just needs to notice equations (3.3) and the following fact
@@
By taking the dual of the last term from the right hand side of above equation, and using equation (3.14), the left hand side of the above equation is equal to
@@
Since both sides of the above equation could be regarded as elements in the ring $\mathbb{C}[\![q]\!][n_{1},n_{2}]$ and we already have proved the above equation for all $n_{1},n_{2}\in\mathbb{Z}_{>0}$, it follows that the equation should still hold true when replacing $n_{1},n_{2}$ by arbitrary complex numbers $t_{1},t_{2}$. Hence, the first equation (1.5) of Theorem 1.3 is thus proved.
@@
which is exactly the first result in Conjecture 6.3 (a) of [1]. Similarly, the second result of Conjecture 6.3 (a) in [1] follows from the special special case $t_{1}=-t_{2}=t$ of equation (1.6).
@@
From the infinite produce formula, it is easy to see that, in above equation, only terms labeled by $q^{n},n\in 2\mathbb{Z}_{\geq 0}$ make non-trivial contribution. Since $(-1)^{n/2}=(-1)^{\binom{n}{2}}$ when $n\in 2\mathbb{Z}_{\geq 0}$, this corollary is obtained by taking the coefficient of $q^{n}$ on both sides of the above equation for an arbitrary fixed nonnegative integer $n$. $\Box$
@@
$\sum_{\lambda\in\mathcal{P}}q^{|\lambda|}sp_{\lambda}\big{(}\bm{p}(\bm{y})\big{)}s_{\lambda}\big{(}\bm{p}^{\prime}(\bm{y}^{\prime})\big{)}=$ $\frac{\prod_{1\leq i<j}(1-q^{2}y^{\prime}_{i}y^{\prime}_{j})}{\prod_{i,j=1}^{\infty}(1-qy^{\prime}_{i}y_{j})(1-qy^{\prime}_{i}y_{j}^{-1})}.$ (4.4)
Actually, Sundaram [18, 19] provided the above equation as an equality for formal functions in the variables $(x_{1},x_{1}^{-1},...,x_{n},x_{n}^{-1})$ for $l(\lambda)\leq n$. We regard it as an identity in the ring $\mathbb{C}[\![q]\!][\bm{p}(\bm{y}),\bm{p}^{\prime}(\bm{y}^{\prime})]$ and it can also be proved by using the method employed in deriving Theorem 4.2.
@@
$\sum_{\lambda\in\mathcal{P}}q^{|\lambda|}\prod_{u\in\lambda}\frac{\big{(}n_{1}+c_{sp}(u)\big{)}\big{(}n_{2}+c(u)\big{)}}{h^{2}(u)}=\prod_{n\geq 1}\frac{(1-q^{2})^{\binom{n_{2}}{2}}}{(1-q)^{2n_{1}n_{2}}}.$ (4.5)
@@ 5. Conflict of interest and data availability statement
Acknowledgements. The author would like to thank Professors Amdeberhan, Andrews, and Ballantine for valuable discussions and their interest in this work. The author would also like to express gratitude to Professors Xiaobo Liu, Jian Zhou, and Xiangyu Zhou for their encouragement. The author is supported by the National Natural Science Foundation of China (No. 12288201) and the China Postdoctoral Science Foundation (No. 2023M743717).
@@
  • [12] D. E. Littlewood. The theory of group characters. Second edition. Oxford University Press, 1990.
@@
  • [23] C. Yang. On two families of Nekrasov-Okounkov type formulas. arXiv:2312.08218, 2023.

### v3

@@ 4.1. Partition function of double symplectic Schur functions
In this subsection, we obtain a formula for the partition function consisting of double symplectic Schur functions as in Theorem 4.2.
Denote the dual operator of $G$ by
$G^{*}:=-\sum_{n=1}^{\infty}\frac{1}{2n}(\alpha_{n}^{2}-\alpha_{2n}).$
We first need the following commutation relation.
@@
Use the variables $(\bm{y})=(y_{i})$ and write $p_{k}$ as $p_{k}(\bm{y})$, the $k$-th power sum of the variables $(\bm{y})$. Then, we have
Lemma 4.3.
The Theorem 4.2 is equivalent to the following
@@
Proof: This is similar to the proof of Lemma 3.2. One just needs to notice equation (3.3) and the following fact
@@
By taking the dual of the last term from the right hand side of the above equation, and using equation (3.21), the left hand side of the above equation is equal to
@@
Since both sides of the above equation could be regarded as elements in the ring $\mathbb{C}[\![q]\!][n_{1},n_{2}]$ and we already have proved the above equation for all $n_{1},n_{2}\in\mathbb{Z}_{>0}$, it follows that the equation should still hold when replacing $n_{1},n_{2}$ by arbitrary complex numbers $t_{1},t_{2}$. Hence, the first equation (1.5) of Theorem 1.3 is thus proved.
@@
which is exactly the first result in Conjecture 6.3 (a) of [1]. Similarly, the second result of Conjecture 6.3 (a) in [1] follows from the special case $t_{1}=-t_{2}=t$ of equation (1.6).
@@
From the infinite product formula, it is easy to see that, in the above equation, only terms labeled by $q^{n},n\in 2\mathbb{Z}_{\geq 0}$ make non-trivial contribution. Since $(-1)^{n/2}=(-1)^{\binom{n}{2}}$ when $n\in 2\mathbb{Z}_{\geq 0}$, this corollary is obtained by taking the coefficient of $q^{n}$ on both sides of the above equation for an arbitrary fixed nonnegative integer $n$. $\Box$
@@
$\sum_{\lambda\in\mathcal{P}}q^{|\lambda|}sp_{\lambda}\big{(}\bm{p}(\bm{y})\big{)}s_{\lambda}\big{(}\bm{p}^{\prime}(\bm{y}^{\prime})\big{)}=$ $\frac{\prod_{1\leq i<j}(1-q^{2}y^{\prime}_{i}y^{\prime}_{j})}{\prod_{i,j=1}^{\infty}(1-qy^{\prime}_{i}y_{j})}.$ (4.4)
Actually, Sundaram [22, 23] provided the above equation as an equality for formal functions in the variables $(x_{1},x_{1}^{-1},...,x_{n},x_{n}^{-1})$ for $l(\lambda)\leq n$. We regard it as an identity in the ring $\mathbb{C}[\![q]\!][\bm{p}(\bm{y}),\bm{p}^{\prime}(\bm{y}^{\prime})]$. For completeness, in the following Lemma 4.5, we provide a proof of the equation (4.4) using the method employed in deriving Theorem 4.2.
@@
$\sum_{\lambda\in\mathcal{P}}q^{|\lambda|}\prod_{u\in\lambda}\frac{\big{(}n_{1}+c_{sp}(u)\big{)}\big{(}n_{2}+c(u)\big{)}}{h^{2}(u)}=\frac{(1-q^{2})^{\binom{n_{2}}{2}}}{(1-q)^{n_{1}n_{2}}}.$ (4.5)
@@
Below, we give a proof of equation (4.4) using the method developed in this paper.
Lemma 4.5.
We have
$\sum_{\lambda\in\mathcal{P}}q^{|\lambda|}sp_{\lambda}\big{(}\bm{p}(\bm{y})\big{)}s_{\lambda}\big{(}\bm{p}^{\prime}(\bm{y}^{\prime})\big{)}=$ $\frac{\prod_{1\leq i<j}(1-q^{2}y^{\prime}_{i}y^{\prime}_{j})}{\prod_{i,j=1}^{\infty}(1-qy^{\prime}_{i}y_{j})}.$ (4.6)
Proof: First, from the Cauchy identity for Schur functions, it is known
$\sum_{\lambda\in\mathcal{P}}s_{\lambda}(\bm{p})\cdot\langle\lambda|=\langle 0|\Gamma_{+}(\bm{p}).$
Then from equations (3.18) and (3.21), we have
$\begin{split}\sum_{\lambda\in\mathcal{P}}q^{|\lambda|}sp_{\lambda}\big{(}\bm{p}^{\prime}(\bm{y}^{\prime})\big{)}s_{\lambda}\big{(}\bm{p}(\bm{y})\big{)}=\langle 0|\Gamma_{+}\big{(}\bm{p}^{\prime}(\bm{y}^{\prime})\big{)}\cdot q^{L_{0}}\cdot\Gamma_{-}\big{(}\bm{p}(\bm{y})\big{)}\cdot\exp(G)|0\rangle.\end{split}$
The right hand side of above equation can be computed using the commutation relations (2.5), (2.6) and (3.12). It is equal to
$\frac{1}{\prod_{i,j=1}^{\infty}(1-qy^{\prime}_{i}y_{j})}\cdot\langle 0|q^{L_{0}}\cdot\Gamma_{-}\big{(}\bm{p}(\bm{y})\big{)}\cdot\exp(G)\cdot\prod_{i=1}^{\infty}\Gamma_{-}(qy_{i}^{\prime})^{-1}\Gamma_{+}(qy_{i}^{\prime})\cdot|0\rangle.$
Then from commutation relations (2.5), (2.6), and Lemma 3.3 again, we can move the operators of the form $\Gamma_{-}(\cdot)$ to the left side. The above equation is equal to
$\frac{\prod_{1\leq i<j}(1-q^{2}y^{\prime}_{i}y^{\prime}_{j})}{\prod_{i,j=1}^{\infty}(1-qy^{\prime}_{i}y_{j})}\cdot\langle 0|\Gamma_{-}\big{(}\bm{p}(q\bm{y})\big{)}\cdot\prod_{i=1}^{\infty}\Gamma_{-}(q^{2}y_{i}^{\prime})^{-1}\cdot q^{L_{0}}\cdot\exp(G)\cdot\prod_{i=1}^{\infty}\Gamma_{+}(qy_{i}^{\prime})\cdot|0\rangle.$
This lemma then follows from the equation (2.9) and $\langle 0|\exp(G)|0\rangle=0$. $\Box$
Remark 4.6.
It is suggested by the anonymous referee that some other specialisations of identities derived in this paper may be interesting. For example, the following specialisation of symplectic Schur functions was considered in [4],
$sp_{\lambda,2n}(s,s^{-1},s^{3},s^{-3},\dots,s^{2n-1},s^{-2n+1})=\prod_{u\in\lambda}\frac{\langle 2n+c_{sp}(u)\rangle_{s}}{\langle h(u)\rangle_{s}},$ (4.7)
where $\langle i\rangle_{s}:=s^{i}-s^{-i}$. This kind of specialisation for Schur functions is known as (see, after some changes of notation, I.3.Example 3 in [15])
$s_{\lambda}(t^{-2m+1},t^{-2m+3},\dots,t^{2m-1},0,\dots)=\prod_{u\in\lambda}\frac{\langle 2m+c(u)\rangle_{t}}{\langle h(u)\rangle_{t}},$ (4.8)
where, similarly, $\langle i\rangle_{t}:=t^{i}-t^{-i}$. Then, applying the specialisations (4.7) and (4.8) to the identity (4.4) gives
$\begin{split}\sum_{\lambda\in\mathcal{P}}&q^{|\lambda|}\prod_{u\in\lambda}\frac{\langle 2n+c_{sp}(u)\rangle_{s}\cdot\langle 2m+c(u)\rangle_{t}}{\langle h(u)\rangle_{s}\cdot\langle h(u)\rangle_{t}}=\frac{\prod_{i=1}^{4m-3}(1-q^{2}t^{2i-4m+2})^{d_{i,m}}}{\prod_{i=1}^{2m}\prod_{j=1}^{2n}(1-qs^{2j-2n-1}t^{2i-2m-1})},\end{split}$ (4.9)
where for $31\leq i\leq 4m-$ 3, define
$d_{i,m}:=\begin{cases}\lfloor(i+1)/2\rfloor,&1\leq i\leq 2m-1,\\ \lfloor(4m-1-i)/2\rfloor,&2m\leq i\leq 4m-3.\end{cases}$
For examples, When setting $n=m=1$, the both sides of equation (4.9) are equal to
$1+\frac{(t^{2}+1)(s^{2}+1)}{st}q$ $+\frac{(s^{4}+s^{2}+1)(t^{4}+t^{2}+1)}{s^{2}t^{2}}q^{2}$
$+\frac{(s^{2}+1)(s^{4}+1)(t^{2}+1)(t^{4}+1)}{s^{3}t^{3}}q^{3}+O(q^{4}).$
When setting $n=2,m=3$, the both sides of equation (4.9) are equal to
$1+\frac{(s^{6}+s^{4}+s^{2}+1)(t^{10}+t^{8}+t^{6}+t^{4}+t^{2}+1)}{s^{3}t^{5}}q+O(q^{2}).$
@@ 5. Conflict of interest and data availability statement
Acknowledgements. The author is grateful to the anonymous referee for many excellent suggestions. The author would like to thank Professors Amdeberhan, Andrews, and Ballantine for their valuable discussions and their interest in this work. The author would also like to express gratitude to Professors Xiaobo Liu, Jian Zhou, and Xiangyu Zhou for their encouragement. The author is supported by the NSFC grants (No. 12288201, 12401079), the China Postdoctoral Science Foundation (No. 2023M743717), and the China National Postdoctoral Program for Innovative Talents (No. BX20240407).
@@
  • [8] F. Garvan, D. Kim, and D. Stanton. Cranks and $t$-cores. Invent. Math., 101(1):1–17, 1990.
@@
  • [12] N. Jing, Z. Li, and D. Wang. Skew symplectic and orthogonal Schur functions. SIGMA Symmetry Integrability Geom. Methods Appl., 20:Paper No. 041, 23, 2024.
@@
  • [14] D. E. Littlewood. The theory of group characters and matrix representations of groups. pages viii+314, 2006. Reprint of the second (1950) edition.
@@
  • [16] W. Miller. Symmetey group and their applciations. Academic Press, New Yourk, 1972.
@@
  • [19] L. Shi, N. Wang, and M. Chen. The orthogonal and symplectic Schur functions, vertex operators and integrable hierarchies. J. Nonlinear Math. Phys., 28(3):292–302, 2021.
@@
  • [27] C. Yang. On two families of Nekrasov–Okounkov type formulas. arXiv:2312.08218, 2023.